OpenStudy (anonymous):

find all values of x in the interval [0,2π] that satisfy the equation, tan(x)-√3=0

OpenStudy (anonymous):

Notice that \[ \tan (\pi/3)= \sqrt 3 \] Hence we have to solve \[ \tan(x) = \tan (\pi/3) \]

OpenStudy (anonymous):

Look at http://www.math.missouri.edu/escgi/mucgi-bin/trigidentities.cgi to see that \[ \tan(x) = \tan (\pi/3) \implies x=\frac \pi 3 + k \pi, \quad k=0,\pm1, \pm2 ,\pm3 \dots\\ x=\frac \pi 3 \\ x=\frac \pi 3 +\pi= \frac {4\pi}3 \] Two solutions between 0 and \( 2 \pi\)